During the course of a​ year, the labor force consists of the same 1,196 people. Of​ these, 20 lack skills that employers desire and hence remain unemployed throughout the year. At the same​ time, every month during the​ year, 26 different people become​ unemployed, and 26 other different people who were unemployed find jobs. There is no cyclical unemployment.

a. Calculate the frictional unemployment rate. %. ​(Enter your response as a percentage rounded to two decimal​ places). 

  (b) Calculate the unemployment rate

-Calculate the frictional unemployment rate -

(d) Calculate the unemployment rate -
Answer #1

Frictional unemployment refers to the type of unemployment in which people leave their jobs to find new jobs. This can be calculated as number of people seeking for job/total number of labour force*100. So according to the information given this can be calculated as:

26/1196*100=2.17%

Unemployment rate refers to the rate of unemployment which includes people who are unemployed, people seeking for job and are unemployent, and , poeple who do not have skills and are unemployed. So total unemployed workers= 20+26+26= 72

Number of unemployed= 72 , total labour force= 1196

Unemployment rate= 72/1196*100=6.02%